Link to report from morning meeting
- 06:03 Dumped fill #3071, integrated lumi 143 pb-1
- Start of TS3
|Sunday 16 September
- 09:30 Fill #3067 dumped by operator. Integrated lumi 165 pb-1.
- 11:21 Start ramp fill #3068.
- 11:36 Dumped at flat top due to beam losses & vacuum IP4. Switch off
- 13:28 Start ramp fill #3069.
- 14:13 Dumped at end of adjust on losses TCP B1.
- 16:21 Start ramp fill #3070.
- 17:01 Dumped end of adjust, carbon copy of previous dump.Change
- 19:20 Start ramp fill #3071.
- 20:08 Stable beams, intial lumi 6.5E33 cm-2s-1.
|Saturday 15 September
- 04:20 all p+ bunches injected.
- 04:35 Dumped on BPMS: pPb did not make it, but important
- Changed LHCb polarity and commission new collision beam
- 06:38 Ramped with 2 x 2 x 6 bunches
- 08:12 Dump by operator
- 10:00 Q20 injection test. Problems with the SPS orbit at
- 12:00 Stop of Q20, revert back to Q26.
- Access for RF and injection kicker. Precycle.
- 14:30 Injection.
- 15:30 Quick access for ATLAS, safety alarm.
- 16:00 Steering lines, reconfigure RF.
- 18:10 Filling for physics.
- 19:50 Stable beams fill #3067,
Luminosity ~6.3E33 cm-2s-1.Some troubles to find LHCb.
- 00:00 Next attempt.(no 4).
- 01:00 Test of OFB at injection.
- 01:30 Ramp and de-squeeze starting. At each stop (90m,
500m) exchange of optics.
- 03:15 At 1 km. Optics measurements.
- 07:00 Dump.
- 08:00 Refill with 3 nominal bunches for highbeta.
- 11:15 Beams colliding at 1 km. Start of collimator and
- 12:45 Dump.
- Access for RF, line 2B1.
- 17:30 pPb MD. Problems to inject 200 ns p+ beam.
- 21:30 Again trip of line 2B1.
- Problems with interlocked BPMS in IP6 and used p+ bunch
- 00:50 Start of loss maps
- 01:26 Stable beams, first time pPb fill #3056
- 06:00 Displace collision point by - 50 cm
- 07:40 Moved collision point to +50 cm
- 09:35 Dump pPb fill #3056. RF access.
- 11:30 Access finished, problems with RF controls.
- 12:30 RF hardware problem back - configure to run with
one klystron less.
- 13:00 Probe beams injected. PBMs need rephasing
- 14:07 SPS mains problem
- 15:00 Inject 3 probes for each beam.
- 15:21 Ramp started
- 15:44 Beams dumped during the first step of squeeze,
collimators energy limits.
- 18:00 Second ramp with 3 probe bunches.
- 19:30 Lost 1/2 B1 at 90 m due to orbit FB - orbit V for
B1 starting to oscillate.
- Squeeze blocked due to a limitation on RTQX circuits 1
and 5. EPC to revert, but requires to switch off PC.
- 20:30 Beams dumped.
- New attempt.
- 22:00 Ramp ok. In the squeeze to 90m ok. Then orbit FB
instability in squeeze to 500m, manually kept the beams. At
500m while investigating the issue, crash of the OFC.
Situation could not be recovered.
- Pumping, removal of vacuum equipment from tunnel in the
- 11:50 pre-cycle started
- 13:00 RF access for klystron problem
- 16:00 Starting injection, Problems with ions, timing
events not sent out correctly.
- 18:30 Filling ions.
- 19:00 Start of ramp. Lost the beam on TCT position
- QPS problems. RF problems.
- 22:52 At 4 TeV with 15 p and 15 Pb bunches
- 00:00 Collisions in all 4 IPs.
|Tuesday 11 Sep
- Overnight leak detection - leak localized on bellow of
WS B1. Attempts to varnish failed.
- Vacuum leak repair ongoing.
Mon 10 Sep
- 06:00 Operator dump, fill #3047, 146pb-1 in 10h36m
- 08:00 Start MD (RF -rephasing in preparation for pPb)
- In the morning, dry tests, then successful tests with beams at 450 GeV, Not
enough time left for tests at 4 TeV.
- 16:00 End of rephasing MD.
- Setting up for p - Pb, Lot's of last minute fixes on SPS and LHC
- 20:48 First Lead injection - right into the correct bucket.
- Setting up RF and ADT loops (gains etc).
- 22:00 B1 dumped. Vacuum valve B1 in IR4.
|08:30 meeting - weekly summary
Inner triplet trip (Markus Zerlauth)